OpenStudy (anonymous):

Solve the following system using substitution or elimination: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Be sure to show all work.

OpenStudy (hba):

@LilyWatson where is the diag

OpenStudy (lgbasallote):

the perimeter is 46 that means x + x + y = 46 it says y is 4 cm longer than x so y = x + 4 sub that x + x + (x+4) = 46

OpenStudy (lgbasallote):

now solve for x and y

OpenStudy (anonymous):

Thank you!